# A Minimalist Optimizer Design for LLM Pretraining

### v1

@@ 5 Conclusion
In this paper, we design a memory-efficient optimizer using a minimalist approach. The proposed algorithm utilizes the building blocks that lead to the success of Adam but further refines them to make it more memory efficient. We motivate each of our construction steps by theoretical or experimental tests. The resulting algorithm, SCALE, achieves superior pretraining performance while only requires 35-45% of Adam memory. This makes the proposed algorithm a strong candidate for large-scale pretraining under memory constraints, as well as a minimalist baseline for benchmarking more sophisticated optimizers.
@@ A.1 Details of memory estimation for 1B and 7B models
Here we compute the memory estimate for both 1B and 7B LLaMA models. We only compute the major parameters, including embedding layers, attention and MLP layers. We follow prior works (Zhao et al., 2024; Han et al., 2024) in estimating the memory using bfloat16 format, where each floating point number occupies 2 bytes.
@@
  • Adafactor: Apart from the parameter states, Adafactor stores a row-wise and column-wise momentum, which is 0.005G memory. In total, Adafactor requires 13.481G memory.
@@
  • Adafactor: Apart from the parameter states, Adafactor stores a row-wise and column-wise momentum, which is 0.002G memory. In total, Adafactor requires 2.68G memory.
@@ A.2 Details of the experiments
We performed wandb sweeps for all methods we tested up to models of size 350M, searching learning rates within {0.00005, 0.0001, 0.0003, 0.0005, 0.001, 0.003, 0.005, 0.01}. For the 1B model, due to resource constraints we manually tune the learning rates using as starting point the optimal learning rate from the 350M sweep. For SCALE, we use learning rate 1e-3 for models sizes 60M, 130M and 350M, 2e-4 for the 1B and 1e-4 for the 7B model. In addition, our reported result for the 1B model uses the same learning rate scaling technique used by Muon (Liu et al., 2025), as we find to result to a minor gain; without it, the achieved perplexity for the 1B model is 14.34. Also, we set the last layer’s momentum parameter $\beta=0.9$, being a common choice for first order momentum. In addition, for all vector parameters we employ the Adam optimizer, following Jordan et al. (2024); Liao et al. (2024). This does not influence the memory usage because the vector parameters are orders of magnitude smaller in size compared to the matrix parameters.
A.3 Curves of Training Iteration versus Perplexity
Figure 4: The evaluation perplexity curves of different methods on LLaMA 1B pretraining. Note that Muon is converging fastest at the beginning stage, while SCALE, Stable-SPAM and APOLLO catch up with Muon in the final stage of training.
A.4 Learning Rate Sensitivity analysis
@@
In Figure 5 we test the performance of our algorithm, SCALE, with different learning rates and compare it with that of Stable-SPAM (Huang et al., 2025a). We observe that both algorithms behave similarly with a reasonable range of learning rates.
A.5 Proofs for Section 2.2
In this section, we conduct the proof for Theorem 2.1. The proof follows Liu et al. (2020). First, we have the following lemmas, which are variations of Liu et al. (2020, Lemma 1 and 2).
@@
for all $l=1,...,L$.
We can take (since $11-\beta_{l}^{t}<$ 1)
@@
$-\frac{3}{4}\eta_{l}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{l,1}\eta_{l}^{2}+4s\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\leq-\frac{\eta_{l}}{2},$ (A.16)
@@
for all $l=1,2,...,L$. Here each term in the right hand side of (A.17) is bounding each term on the left hand side of (A.16).
@@
Now by telescoping sum of (A.22) for $t=1,...,T$, also notice that $1/\eta=\Theta(\sqrt{\gamma T}L\gamma/\delta^{2})$ as $T$ grows, we get our final result of

### v2

@@ 5 Conclusion
In this paper, we design a memory-efficient optimizer using a minimalist approach. The proposed algorithm utilizes the building blocks that lead to the success of Adam but further refines them to make it more memory efficient. We motivate each of our construction steps by theoretical or experimental evidences. The resulting algorithm, SCALE, achieves superior performance to existing memory-efficient optimizers for LLM pretraining and matches Adam while only requiring 35-45% of memory. This makes the proposed algorithm a strong candidate for large-scale pretraining under memory constraints, as well as a minimalist baseline for benchmarking more sophisticated optimizers.
@@ A.1 Details of memory estimation for 1B and 7B models
Here we compute the memory estimate for both 1B and 7B LLaMA models. We only compute the major parameters, including embedding layers, attention and MLP layers. We follow prior works (zhao2024galore; han2024sltrain) in estimating the memory using bfloat16 format, where each floating point number occupies 2 bytes.
@@ A.2 Details of the experiments
For all LLaMA experiments, we follow zhao2024galore and set the sequence length to 256 and the batch size to 512, train using BF16 format and report evaluation perplexity as our metric. We also use a cosine learning rate with linear warm-up for the first 10% of the iterations. For low-rank optimizers (GaLore, Fira and APOLLO) we follow their suggested hyperparameters (including the rank) but tune the learning rates. For Muon we follow the implementation from liu2025muon.
@@
We performed wandb sweeps for all methods we tested up to models of size 350M, searching learning rates within {0.00005, 0.0001, 0.0003, 0.0005, 0.001, 0.003, 0.005, 0.01}. For the 1B model, due to resource constraints we manually tune the learning rates using as starting point the optimal learning rate from the 350M sweep. For SCALE, we use learning rate 1e-3 for models sizes 60M, 130M and 350M, 2e-4 for the 1B and 1e-4 for the 7B model. In addition, our reported result for the 1B model uses the same learning rate scaling technique used by Muon (liu2025muon). Also, we set the last layer’s momentum parameter $\beta=0.9$, being a common choice for first order momentum. In addition, for all vector parameters we employ the Adam optimizer, following jordan2024muon; liao2024galore. This does not influence the memory usage because the vector parameters are orders of magnitude smaller in size compared to the matrix parameters.
A.3 Training Throughput Comparison of Different Methods
We conduct a throughput (tokens/sec) comparison of the different optimizers for training LLaMA 1B on 4xH100 GPUs using the same settings as specified in A.2. Table 7 shows that our method is about $18.5\%$ faster than singular-value (NS) based normalization methods (Muon/SWAN) and about $8\%$ faster than GaLore/Fira which project the Adam states to low-rank using SVD. Moreover, it achieves similar throughput to Adam/Stable-SPAM and APOLLO(-Mini).
Method Throughput (tokens/sec)
Adam 45019
Adam (Stable-SPAM) 44960
Singular-value-based (Muon/SWAN) 37748
GaLore 41267
Fira 41285
APOLLO 44193
APOLLO-Mini 44567
SCALE 44728
Table 7: Throughput comparison of different methods for training LLaMA 1B on 4xH100 GPUs. $^{\star}:$ Due to code unavailability we cannot directly test the throughput of SWAN, instead we report the throughput using the Newton-Schulz (NS) approximation of Singular-value normalization (see jordan2024muon for details).
A.4 Pretraining Results on Additional LLM Architectures
For our main experiments we choose the LLaMA family of models as it is commonly used to benchmark memory efficient optimizers for LLM pretraining (such as in GaLore (zhao2024galore), Fira (chen2024fira), SWAN (ma2024swan) and APOLLO (zhu2025apollo)), which is the scope of our paper. However, to provide further evidence of the generality of our algorithm among LLM architectures, we conducted additional pretraining experiments with Qwen2-500M (yang2024qwen2technicalreport), GPT2-Medium (355M) (radford2019language) and Gemma-2B (team2024gemma). We followed the same experimental settings as in our LLaMA experiments, described in A.2. We provide the results bellow:
Model Qwen2-500M GPT2-M (355M)
Tokens 10B 7.8B
Adam 17.61 (2.96G) 20.73 (2.13G)
Adam (Stable-SPAM) 15.91 (2.96G) 18.90 (2.13G)
Muon 16.03 (1.98G) 19.61 (1.42G)
GaLore 18.22 (1.94G) 23.66 (1.22G)
Fira 15.94 (1.94G) 19.41 (1.22G)
APOLLO 16.04 (1.94G) 19.30 (1.22G)
APOLLO-Mini 16.17 (1.53G) 19.99 (0.92G)
SCALE 15.57 (1.26G) 19.00 (0.81G)
Table 8: Experiment results from pretraining Qwen and GPT2 models on the C4 dataset.
From Table 8 we observe that our method still achieves Adam-like performance while using significantly less memory and continues to outperform SOTA memory-efficient methods. We believe that our simple, yet systematic, optimizer design makes our method more robust and less likely to overfit to specific LLM architectures.
Model Gemma-2B
Tokens 40B
APOLLO 11.90 (9.09G)
SCALE 11.80 (6.06G)
Table 9: Experiment results from pretraining Gemma-2B on the C4 dataset. Due to resource and time constrains we limit our comparison to APOLLO, being the strongest baseline.
From table 9 we see that SCALE achieves a lower perplexity with an even smaller memory footprint than APOLLO, which further showcases its SOTA memory-efficient pretraining performance.
A.5 Results on Overtraining Regime
Chinchilla ratio $1\times$ $2\times$ $4\times$
Adam (2.21G) 18.77 17.60 17.21
Adam (Stable-SPAM) (2.21G) 16.80 15.85 15.11
Muon (1.47G) 16.70 15.81 15.18
GaLore (1.59G) 19.37 18.40 17.81
Fira (1.59G) 16.82 15.82 15.31
APOLLO (1.59G) 16.75 15.76 15.06
APOLLO-Mini (1.00G) 17.11 16.02 15.21
SCALE (0.80G) 16.32 15.33 14.77
Table 10: Results from training the 350M LLaMA model on C4 using different token budgets. Chinchilla ratio $1\times$ corresponds to (roughly) our default token budget for the given model, i.e., 7.8B tokens (following zhao2024galore), $2\times$ to 15.6B tokens and $4\times$ to 31.2B tokens. Our method maintains its SOTA memory-efficient pretraining performance among the different training budgets.
A.6 Fine-tuning Results
First, we want to emphasize that our optimizer design is focused on memory-efficient LLM pretraining. As pretraining is by far the most computationally expensive part of the LLM training process (for example pretraining LLaMA2-70B took about 1.72 million A100 GPU hours (touvron2023llama)), it is commonly believed that an efficient SOTA pretraining optimizer is of great importance.
Instead, for fine-tuning, existing Parameter-efficient fine-tuning (PEFT) methods already reduce the memory and compute requirements substantially while giving performance competitive to full fine-tuning. Therefore we don’t expect our optimizer to replace such approaches. However, to give further evidence of its generality, we provide some preliminary results on GLUE benchmarks with a pretrained RoBERTa-base (liu2019roberta) which we fine-tuned with our method and Adam (full fine-tuning), following the setup of zhao2024galore. We conduct a learning rate search for both methods and report the best epoch results (higher the better).
Method / Benchmark RTE CoLA MRPC STS-B SST-2 QNLI Avg
Adam           (0.75G) 79.53 63.37 93.01 91.26 94.26 92.67 85.68
SCALE           (0.33G) 80.46 63.02 92.53 91.23 93.58 92.23 85.51
Table 11: Results from fine-tuning RoBERTa-base on different GLUE benchmarks (wang2018glue). Memory consumption is indicated in the parenthesis.
We observe that our method, despite being designed for pretraining, can give results comparable to Adam for fine-tuning tasks while using over 2 times less memory (included in parenthesis). Again, we want to point out that our optimizer is aimed at memory-efficient pretraining where its advantages are most relevant. As a future work, it might be worth investigating combining SCALE with PEFT methods, although this is out of the scope of this paper.
A.7 Learning Rate Sensitivity Analysis
@@
In Figure 5 we test the performance of our algorithm, SCALE, with different learning rates and compare it with that of Adam (Stable-SPAM version) (huang2025stablespam). We observe that both algorithms behave similarly with a reasonable range of learning rates.
A.8 Curves of Training Iteration versus Perplexity
Figure 6: The evaluation perplexity curves of different methods on LLaMA 1B pretraining. Note that Muon is converging the fastest at the beginning stage, while SCALE, Adam (Stable-SPAM) and APOLLO-Mini catch up in the final stage of training.
A.9 Discussion on why Column-wise matters for the Last Layer
(a) Iteration 1,000.
(b) Iteration 10,000.
Figure 7: We present the column-norm values of the last layer’s gradient matrix at iterations 1,000 and 10,000 during the training of LLaMA 130M. The last layer of LLaMA 130M has gradient dimensions $d_{model}\times|V|=768\times 32,000$, which means that there are $|V|=32,000$ columns (and therefore $|V|$ corresponding column-norms). For clarity we limit the x axis to the first 1000 columns. We also note that due to the design of the SentencePiece tokenizer (kudo2018sentencepiece) used, lower token ids (and therefore lower last-layer column-ids) generally correspond to more frequent tokens (such as “_the”, “_and”, “_to”, etc., as we observed). Therefore, from the above figures it can be seen that more frequent tokens have much larger column-norms, potentially leading to imbalanced learning. We hypothesize that the effectiveness of column-wise is related to mitigating this phenomenon by normalizing all the column norms to the same level and thus enabling more balanced training dynamics.
In this section we further investigate the importance of column-wise normalization for the last layer, and also ablate several “mixed” normalization schemes, where more than one type of normalization is used depending on the layer.
Method Perplexity
1. SCALE (all column, mmt-last) 22.57
2. column-last, row-rest, mmt-last 23.27
3. row-first, column-rest, mmt-last 22.94
4. norm along larger dim, mmt-last 23.52
5. row-last, column-rest, mmt-last 28.83
Table 12: Ablation results from training the 130M LLaMA model on C4 using “mixed” normalization schemes compared to SCALE.
All results of Table 12 are from pretraining the LLaMA 130M model using momentum only for the last layer (mmt-last, as in SCALE). Methods 1-4 all do column-wise normalization for the last layer; only method 5 (row-last, column-rest, mmt-last) does row-wise for the last layer (row-last). Note that “column-rest” or “row-rest” means we do column- or row-wise normalizations for all other layers we don’t explicitly state.
We can clearly observe from the above results that row-last significantly degrades the performance. In addition, we can see that excluding the last layer, the rest of the layers are less influenced by the normalization direction. However, a uniform normalization approach as in SCALE (all column-wise) still outperforms other “mixed” normalization approaches.
To explain the above results, our intuition suggests that the last layer in LLMs is somehow “special”. Indeed, it is the linear projection layer (often called Language modeling head) that maps the last hidden state (of dimension $d_{model}$) to the vocabulary logits (of dimension $|V|$ being the vocabulary size). Its gradient matrix shape is $d_{model}\times|V|$, i.e., it has $|V|$ columns, same as its weight matrix shape. Therefore, each gradient column of the last layer has a “physical meaning”, i.e., it corresponds to one of the $|V|$ vocabulary tokens.
We have observed that for certain (few) tokens, their corresponding last-layer column gradients have much larger norms compared to the rest. We have also observed that such columns correspond to more frequent tokens (see Figure 7 details). We hypothesize that such last-layer column-norm differences can potentially lead to uneven learning (i.e., more rare tokens receive smaller gradients and therefore are not “learned” by the model) or even training divergence (due to the large gradients corresponding to frequent tokens). In connection to this, kunstner2024heavy, aiming at explaining the large Adam-SGD gap in LLM training, observed that SGD is able to make much less progress on learning low-frequency classes compared to Adam. Based on those insights, we conclude that by normalizing the columns of the last layer our method obtains more stable training dynamics and a more “even” token learning behavior, even without using full per-parameter adaptivity as in Adam.
Although still at a preliminary stage, we believe that our above insights for the importance of column-wise normalization for the last layer can provide a deeper understanding of the normalization component and (at least to some degree) explain why SCALE is able to perform comparable to Adam for LLM pretraining.
A.10 Proofs for Section 2.2
In this section, we conduct the proof for Theorem 2.1. The proof follows liu2020improved. First, we have the following lemmas, which are variations of liu2020improved.
@@ A.9 Discussion on why Column-wise matters for the Last Layer
for all l=1,,Ll=1,...,L.
We can take (since $1-\beta_{l}^{t}<1$)
@@
$\frac{3}{4}\eta_{l}+\frac{\gamma\eta_{l}^{2}}{2}+4c_{l,1}\eta_{l}^{2}+4s\eta_{l}^{2}(\frac{\beta_{l}}{1-\beta_{l}})^{2}\gamma^{2}(1-\beta_{l}^{t-1})^{2}\leq-\frac{\eta_{l}}{2},$ (A.16)
@@
for all l=1,2,,Ll=1,2,...,L. Here each term in the right hand side of (A.17) is bounding each term on the left hand side of (A.16).
@@
Now by telescoping sum of (A.22) for t=1,,Tt=1,...,T, also notice that $1/\eta=\Theta(\sqrt{\gamma T}L\gamma/\delta^{2})$ as $T$ grows, we get our final result of